Tell whether the equation is linear. If so, write the equation in standard form and give the values of A, B, and C.
2x = y/3 - 4

Answer by mananth(16949)

2x = y/3 - 4
2x-y/3 +4 =0
...
A=2,
B=-1/3
C=4